# The New Science of Unidentified Aerospace-Undersea Phenomena (UAP)

### v1

@@ 4.1.6 VASCO
4.2 Private Programs
@@ 4.2.2 Eye on the Sky
The Eye on the Sky platform is a fully instrumented mobile sensor developed by John and Gerald Tedesco, enabling the observation, detection, tracking, data collection, and identification of unidentified objects in our skies and along our coastline. This state-of-the-art experimental expeditionary vehicle, a fully instrumented high-tech mobile surveillance system, offered a centralized location and system integration for environmental surveys and safety assurance. As engineers in this study, the Tedesco’s deployed two iterations of a cutting-edge proto-type mobile laboratory-observatory vehicle. Nightcrawler-1, a robust ‘bottom-up construction design, was used for the first three months of the study. Nightcrawler-2, an RV with a robust ’top-down’ modification design and a retrofitted instrument payload, was used for the subsequent seven months of scientific field study due to its superior speed and mobility. It remains the current mobile system and is in continuous field use.
@@ 5.2 National Aviation Reporting Center on Anomalous Phenomena (NARCAP)
The National Aviation Reporting Center on Anomalous Phenomena (NARCAP) was established late in 1999 by Richard F. Haines (Chief Scientist), (see A.4) and Ted Roe (Executive Director) to deal with the fact that commercial pilots were not reporting the aerial phenomena they had seen, of highly unusual and unexplainable characteristics, to their management. At the time, it was likely that military pilots were doing much the same.
@@
NARCAP called together an advisory group made up of two notable physicists, several engineers and pilots, a meteorologist, an air traffic controller, a human factors specialist, and several others. They provided invaluable advice on technical and policy issues. See Section A.16 in this report regarding Jacques Vallée who also served on our Advisory Board.
How could UAP flying near an airplane possibly be dangerous? First, if the UAP came close enough, the pilot might make a rapid avoidance maneuver (in addition to its various consequences to passengers and subsequent responses by air traffic controllers on the ground). This has happened! The flight crews could then become so fascinated by their presence that they interfere with their continuous monitoring of their flight instruments. This has also happened!
Disturbances of various cockpit instruments by electro-magnetic energy (seemingly) emanating from the UAP could impact flight safety. This has also definitely happened. Having reliable E-M data may contribute to a fuller understanding of the physics related to UAP.
@@ 5.4 The Sentinel Center
As a result, SIGMA2 studies of UAP cases are conducted scientifically by relying on physical data like radar, IR/electro-optical or electromagnetic observables, with the goal of understanding the physics related to the UAP observables, including the effects induced on human beings. For this reason, the efforts made by SIGMA2 are interdisciplinary in nature, simultaneously involving physicists employing optical, infrared, and electromagnetic sensors to study the observables of UAP, but also involving psychologists studying the effects induced on human beings, which include the psychological effects of UAP on people.
Many reports and documents are published in the 3AF-SIGMA2 media repository [304] among them the 3AF-SIGMA2 Progress report 2021 which includes a list of cases analysis and a tour of the physical features of UAP [146]. 3AF-SIGMA2 also organizes webinars on physical observables of UAP [305].
@@ 5.6 The Society for UAP Studies (SUAPS)
The Society for UAP Studies (https://www.societyforuapstudies.org/) is a 501(c)(3) nonprofit organization, which was organized in September 2022 to play a vital role in creating the necessary academic foundations for a stable, sustainable transgenerational scholarly field focused on the many aspects of UAP. The Society takes as its primary mission to promote and actively aid in the development of UAP studies as a proper area of scholarly research.
@@ 5.8 UAP Check
6 UAP and Nuclear Weapons
One of the most disturbing and important aspects of UAP is the fact that they have been observed for decades to be active around nuclear power plants and nuclear weapon sites. The most shocking events involve UAPs hovering over nuclear ICBM missile silos and shutting them down. The most famous of these events are the 1967 incursions at the nuclear weapon sites near Malmstrom Air Force Base (AFB), in Great Falls, Montana, which resulted in the shutdown of almost twenty nuclear missiles [238, 232, 239]. Even more disturbing, journalist George Knapp testified before the US Congress [26] that Russian Col. Boris Sokolov shared information about a similar event that took place at a Russian ICBM base in Ukraine where
This topic, while important to mention, falls outside the scope of this paper on scientific efforts to study UAP. We direct the interested reader to the work of Hastings [232], as well as Robert Salas and James Klotz [238, 239]. In addition, there is an excellent series of studies performed by members of the Scientific Coalition for UAP Studies (SCU), Larry J. Hancock, Ian M. Porritt, Sean Grosvenor, Larry Cates, and Ike Okafor, dealing with the relationship between UAP and the 1945-1975 United States Atomic Warfare Complex [233, 240], and the US military in general [241].
The Hancock et al. SCU study considered three classes of atomic warfare complex sites: 1) radioactive materials production plants, 2) atomic weapons assembly facilities, and 3) atomic weapons stockpile sites. They found that [233]
The SCU findings concluded that these UAPs appeared to be involved in intelligent and focused activity, which is possibly best described as surveillance. This is an important fact to recognize and appreciate, as the sciences are usually focused on studying passive physical phenomena—not phenomena exhibiting intelligence and intention.
7 Transmedium Travel and Water
To use a colloquialism, it does not take a rocket scientist (or an aerospace engineer, for that matter) to recognize and appreciate that craft with effective and capable transmedium abilities are quite beyond human competency. A seaplane is both a very poor airplane and a very poor boat! For this reason, the observed transmedium behaviors of UAP should perhaps give us the most pause.
7.1 Historical Significance of USO Encounters
When we explore the old encounters, we find many intriguing events that never receive the attention they appear to deserve. The following event from 1887 occurred near Cape Race, off the coast of Newfoundland, Canada [244][245, pp. 442–443][246, pp. 167–168][6]. It reads as follows:
[…]
This area was in the Niger Protectorate and Gabon, about 250 miles inland from the West African coast. Kingsley wrote that late at night, she went out alone to bathe and canoe on the lake waters. Then she saw something extraordinary: [249, p. 254]
Down through the forest on the lake bank opposite came a violet ball the size of a small orange. When it reached the sand beach, it hovered along it to and fro close to the ground. In a few minutes another ball of similarly colored light came towards it from behind one of the islets, and the two wavered to and fro over the beach, sometimes circling round each other. I made off toward them in the canoe, thinking — as I still do — they were some brand new kind of luminous insect. When I got onto their beach, one of them went off into the bushes and the other away over the water. I followed in the canoe, for the water here is very deep and, when I almost thought I had got it, it went down into the water and I could see it glowing as it sunk until it vanished in the depths.
7.2 USOs: Unidentified Submerged Objects
[…]
  • objects that are observed entering or falling into the water;
  • objects that are observed leaving the water and rising into the air;
[…]
@@ 8 Social Sciences
An analogy with climate change helps to clarify this point. Beliefs about climate change have significant social consequences that merit study, but climate change itself has tangible societal effects that extend beyond individual or collective beliefs. If researchers indefinitely suspend their judgment on the reality of climate change, they miss the opportunity to study its direct impact on social structures and practices. Similarly, by suspending judgment on the existence of UAP, social scientists limit the range of questions they can ask witnesses and may overlook the potential significance of patterns in testimonies. For example, when witnesses across decades and regions report objects that show intelligent control, incredible speeds, non-aerodynamic shapes and maneuvers, and near-complete silence, a social theorist might focus on how these accounts spread as sociocultural memes and express long-standing cultural anxieties [319]. This focus risks ignoring what these reports might reveal about the underlying phenomenology of whatever is controlling the UAP. As a result, this approach can distort our understanding of witness testimonies by reducing them solely to historical and sociocultural phenomena.
@@ 9 The Scientific Methodology and Best Practices for Collecting UAP Data
In this section we summarize the scientific methodology that has been found to be useful, as well as best practices, to detect, monitor and study UAP. Several papers have been published that present useful summaries. Most notable are the papers by Ailleris [63, 64], the Galileo Project [272, 275, 270, 274, 62], the Hessdalen Project [259, 256, 332, 257, 258], the UAlbany-UAPx Collaboration in this volume [61], the UAP Tracker Project [292, 291], the Tedesco’s Eye On The Sky Project [290, 289], and the numerous IFEX projects, such as SKY-CAM, SONATE-2, KI-SENS, and VaMEx3-MarsSymphony, which are summarized in Sec. 4.1.3.
@@ 9.3.1 Light-Based Imagery
It is extremely important to obtain information about the light beyond the images. Spectra can provide information about the elements or molecules involved in the light production, which, in addition to informing us about the kinds of lights being used, can also provide information about the UAP’s lift and propulsion mechanisms. Obtaining imagery using a polarizer can provide information about the polarized light environment surrounding the UAP. The axis of polarized light can be rotated by strong magnetic fields via a phenomenon known as the Faraday effect [337], Given that the scattered light in the sky is generally polarized, one can use the rotation of the axis of polarization to estimate magnetic field strengths around UAP. Lastly, it may be beneficial to take optical photographs using several different exposure times. This can help to provide valuable information about the speed and trajectory of a UAP in three dimensions, especially if multiple spatially separated cameras are used for triangulation.
@@ Acknowledgements
The authors express their gratitude to Peter Sturrock, who passed away while this paper was being written. As a result, Sturrock was never completely aware of the influence of his efforts and how they have changed the landscape in the context of this UFO phenomenon. Similarly, there are a number of UFO researchers who could have been mentioned along with their colleagues in the appendix. In the end, both space and time became constraining factors, but we wish to emphasize that this in no way lessens our appreciation of their contributions. In this spirit, the authors express their gratitude to Jan Aldrich and Richard Hall.
@@
Last, while it is perhaps awkward to thank one’s coauthors, KHK expresses his gratitude to them for their efforts in studying UAP (some of whom have been doing this his whole life, and then some!), for inspiring him to take on this exciting, albeit daunting, and at times, intimidating, challenge, for assisting in writing this altogether enormous and comprehensive paper, and for several years of friendship and support through what have been the most challenging and exciting years of his life. Most importantly, KHK thanks his family and especially his wife, Emily, for their support.
@@ Appendix A Prominent Past Efforts and Individuals
Four years ago, the director of Harvard College Observatory, the Administrator of NASA, the Director of the CIA and the Archbishop of Washington stood at the National Cathedral to announce that the enigma represented by UAP (UFO) reports remained an unsolved mystery. Following such statements by government and military authorities about the reality of unexplained sightings of unexplained objects in the Earth’s environment, the public and the scientific community have begun to revise their negative impressions of the subject. Naturally, they now seek reliable information about its nature. For this reason, reconstructing and presenting the history of research on the topic of UAP has become critical.
The range of phenomena known as UAP (formerly referred to as UFOs) was never accepted as a subject of interest in the Hard Sciences. Even quality studies into the material nature and behavior of the reported objects have been consistently discouraged by major scientific magazines—journals such as Science, Nature or even Scientific American. Their pages continue to be denied for such studies, even when submitted by established scientists with credentials and a history of publication in other areas.
Much UAP literature of various quality is now found in the form of student theses, conference transcripts, and papers about witness psychology, reporting statistics, medical symptoms, and sociological interpretations of sightings, but not in formal papers about physics, engineering, or site investigation of actual cases. As for computer compilations of the many thousands of reported cases, few sources are publicly available and their reliability is debatable. Such documentation does exist, however, in the form of research studies dating back to the 1960s, but it is easy to miss because it had to be published in rare amateur magazines that have only survived in obscure Internet archives or in the confusion of online depositories. An important exception is the professionally curated and edited magazine called Flying Saucer Review, which was published in England from 1955 to 2000. The researchers who submitted their results to such an obscure publication were particularly dedicated and passionate for the subject (and no, they did not always agree).
In this section, we summarize the research efforts and works from a number of prominent PhD scientists and academics who studied UFOs relatively early on. We explicitly mentioned and described Flying Saucer Review because it is one of the only publications to which many of these individuals were able to submit their works.
@@ A.1 David Akers, Yakima Valley, Washington USA
David Akers coordinated with local residents to study anomalous light phenomena on the Yakima Indian Reservation in the vicinity of Toppenish, WA, from 1972 to the early 2000s. This activity started in 1971 and spiked in the late summer of 1972, consisting largely of nocturnal lights very near the ground or at low altitude. The initial phase of investigation lasted from August 20 to August 31, 1972, during an intense wave of sightings; many of the most interesting results derived from this period. In this interval, nocturnal lights were reported on daily basis in the towns of Union Gap, Harrah, Zillah, Toppenish, Wapato, and Mabton. Akers used optical SLR cameras (VIS-NIR), a compass spin detector, infrared photometer, microphone and casette recorder, and energetic particle detector (Geiger-Muller counter) (see [345] Appendix B). Leveraging community reports, Akers deployed his instruments where sightings were most recently reported (a ‘‘hit and run’’ approach) and was successful witnessing and recording events in this way. The activity subsided dramatically after the first two days of the study.
@@
Akers acquired several photographs of what he described as ‘‘large, flickering orange objects’’ [345], although no quantitative size determination was possible. One object viewed through binoculars was described as a ‘‘sharply defined, brownish orange disc with a mottled surface’’. This object appeared to be flat, and illuminated the ground with an orange glow. The appearance and motion of the light phenomena were highly distinctive and easily distinguished from conventional aircraft. On multiple occasions, attempts to acquire photographs were foiled when illumination ceased abruptly just as a camera was trained on the object. The most striking example of this occurred at relatively close range, while an object hovered directly overhead [345].
@@
On at least one occasion, a light was recorded in a background exposure, captured when no lights were visible to the naked eye, suggesting near-infrared emission, or visible band emission with a pulse duration and repetition rate below the perception threshold of human sight. Akers also noted the possibility that the objects may have been moving in a discrete or noncontinuous fashion, as suggested by the observed flickering (when observed by eye) and discontinuities in long-exposure photographs; this was not definitively distinguished from ordinary flashing and displacement while the objects are nonluminous. During the initial 11-day field campaign, none of the other instruments succeeded in recording the anomalous events. On one occasion, Akers and a colleague experienced over a minute of extremely bright and disorienting light flashes in their peripheral vision, later tentatively ascribed to strong transient magnetic field fluctuations later observed at the same location [345].
A.2 Dr. Erol A. Faruk
@@
Dr. Erol A. Faruk was born in London, UK in 1951. While undertaking postdoctoral organic chemical research at Nottingham University in England he requested and received, through CUFOS, soil material which had reportedly been subjected to physical and chemical changes by an illuminated UFO observed hovering just above the ground. This event [346] occurred in November, 1971 on a farm at Delphos, Kansas, USA and left an eight-foot diameter ring of soil whose surface was found to be glowing brightly once the illuminated UFO had left the area. Subsequent field investigations showed that the ring soil had now become impervious to water down to a depth of 14 inches, such that when snow had fallen later there was no absorption by the ring, leaving only a snow-covered shape while all around was dark muddy soil.
@@
After receiving the material Faruk approached his academic supervisor (Dr B. W. Bycroft) who, after being described the event and associated glow, became curious enough to refer him to the resident chemiluminescence expert (Dr Frank Palmer) who became similarly intrigued by the case. Together they enabled Faruk to analyze the material while Palmer undertook recording of fluorescence spectra on samples taken as the investigation proceeded. Faruk discovered that the ring soil was impregnated by an air-sensitive organic compound that on chromatographic purification revealed it to have all the attributes necessary to generate light, a process known as oxidative chemiluminescence. He was able to use the soil data to propose a viable theory as to how the ring had actually formed, essentially confirming the witnesses’ description of the nearby presence of a hovering aerial device of unknown technology and origin [347, 348].
A.3 Dr. Pierre Guérin, Institut d’Astrophysique
@@
A.4 Dr. Richard F. Haines, NASA-Ames Research Center
@@
A.5 Paul R. Hill, NASA Langley Research Center
Paul R. Hill (1909-1990) was a distinguished American aerodynamicist who made significant contributions to mid-20th-century aerospace research. He obtained in his BS in mechanical engineering from the University of California, Berkeley in 1936, and joined the National Advisory Committee of Aeronautics (NACA), the predecessor to NASA, in 1939. He worked on various NASA aerospace projects, including aircraft design, space missions, hypersonic propulsion, and the Apollo moon missions. He was one of NASA’s chief scientists and was awarded NASA’s Exceptional Service Medal in 1969. The historian James Hansen describes Hill as one of two ‘‘key members of Langley’s early space station research.’’ [374, pp. 275–277], and as the specialist on configuration and propulsion on the HYWARDS (HYpersonic Weapon And R and D System) Program [375, p. 368].

### v2

@@ 7.1.6 VASCO
7.2 Private and Citizen Science Research Efforts
@@ 7.2.2 Eye on the Sky
The Eye on the Sky platform is a fully instrumented mobile sensor developed by John and Gerald Tedesco, enabling the observation, detection, tracking, data collection, and identification of unidentified objects in our skies and along our coastline. This state-of-the-art experimental expeditionary vehicle, a fully instrumented high-tech mobile surveillance system, offered a centralized location and system integration for environmental surveys and safety assurance. As engineers in this study, the Tedescos deployed two iterations of a cutting-edge proto-type mobile laboratory-observatory vehicle. Nightcrawler-1, a robust ‘bottom-up construction design, was used for the first three months of the study. Nightcrawler-2, an RV with a robust ’top-down’ modification design and a retrofitted instrument payload, was used for the subsequent seven months of scientific field study due to its superior speed and mobility. It remains the current mobile system and is in continuous field use.
@@ 8.2 National Aviation Reporting Center on Anomalous Phenomena (NARCAP)
The National Aviation Reporting Center on Anomalous Phenomena (NARCAP) was established late in 1999 by Richard F. Haines (Chief Scientist), (see A.5) and Ted Roe (Executive Director) to deal with the fact that commercial pilots were not reporting the aerial phenomena they had seen, of highly unusual and unexplainable characteristics, to their management. At the time, it was likely that military pilots were doing much the same.
@@
NARCAP called together an advisory group made up of two notable physicists, several engineers and pilots, a meteorologist, an air traffic controller, a human factors specialist, and several others. They provided invaluable advice on technical and policy issues. See Section A.17 in this report regarding Jacques Vallée who also served on their Advisory Board.
How could UAP flying near an airplane possibly be dangerous? First, if the UAP came close enough, the pilot might make a rapid avoidance maneuver (in addition to its various consequences to passengers and subsequent responses by air traffic controllers on the ground). The flight crews could then become so fascinated by their presence that they interfere with their continuous monitoring of their flight instruments.
Disturbances of various cockpit instruments by electro-magnetic energy (seemingly) emanating from the UAP could impact flight safety [353]. This has also definitely happened. Having reliable E-M data may contribute to a fuller understanding of the physics related to UAP.
@@ 8.4 The Sentinel Center
As a result, SIGMA2 studies of UAP cases are conducted scientifically by relying on physical data like radar, IR/electro-optical or electromagnetic observables, with the goal of understanding the physics related to the UAP observables, including the effects induced on human beings. For this reason, the efforts of SIGMA2 are interdisciplinary in nature, simultaneously involving physicists employing optical, infrared, and electromagnetic sensors to study the observables of UAP, but also involving psychologists studying the effects induced on humans, including the psychological effects of UAP on people.
Many reports and documents are published in the 3AF-SIGMA2 media repository [361] among them the 3AF-SIGMA2 Progress Report 2021, which includes a list of cases analysis and a tour of the physical features of UAP [149]. 3AF-SIGMA2 also organizes webinars on physical observables of UAP [362].
@@ 8.6 The Society for UAP Studies (SUAPS)
The Society for UAP Studies (https://www.societyforuapstudies.org/) is a 501(c)(3) nonprofit organization, which was organized in September 2022 to play a vital role in creating the necessary academic foundations for a stable and sustainable transgenerational scholarly field focused on the many aspects of UAP. The Society takes as its primary mission to promote and actively aid in the development of UAP studies as a proper area of scholarly research.
@@ 9 Social Sciences
An analogy with climate change helps to clarify this point. Beliefs about climate change have significant social (and existential) consequences that merit study, but climate change itself has tangible societal effects that extend beyond individual or collective beliefs. If researchers indefinitely suspend their judgment on the reality of climate change, they miss the opportunity to study its direct impact on social structures and practices. Similarly, by suspending judgment on the existence of UAP, social scientists limit the range of questions they can ask witnesses and may overlook the potential significance of patterns in testimonies. For example, when witnesses across decades and regions report objects that show intelligent control, incredible speeds, non-aerodynamic shapes and maneuvers, and near-complete silence, a social theorist might focus on how these accounts spread as sociocultural memes and express long-standing cultural anxieties [376]. This focus risks ignoring what these reports might reveal about the underlying phenomenology of whatever is controlling the UAP. As a result, this approach can distort our understanding of witness testimonies by reducing them solely to historical and sociocultural phenomena.
@@ 10 The Scientific Methodology and Best Practices for Collecting UAP Data
In this section we summarize the scientific methodology that has been found to be useful, as well as best practices, to detect, monitor and study UAP. Several papers have been published that present useful summaries. Most notable are the papers by Ailleris [63, 64], the Galileo Project [329, 332, 327, 331, 62], the Hessdalen Project [316, 313, 389, 314, 315], the UAlbany-UAPx Collaboration in this volume [61], the UAP Tracker Project [349, 348], the Tedescos’ Eye On The Sky Project [347, 346], and the numerous IFEX projects, such as SKY-CAM, SONATE-2, KI-SENS, and VaMEx3-MarsSymphony, which are summarized in Sec. 7.1.3.
@@ 10.3.1 Light-Based Imagery
It is extremely important to obtain information about the light beyond the images. Spectra can provide information about the elements or molecules involved in the light production, which, in addition to informing us about the kinds of lights being used, can also provide information about the UAP’s lift and propulsion mechanisms. Obtaining imagery using a polarizer can provide information about the polarized light environment surrounding the UAP. The axis of polarized light can be rotated by strong magnetic fields via a phenomenon known as the Faraday effect [394], Given that the scattered light in the sky is generally polarized, one can use the rotation of the axis of polarization to estimate magnetic field strengths around UAP.
@@ Acknowledgements
The authors express their gratitude to Peter Sturrock, who passed away while this paper was being written, and whose influence on UAP studies cannot be overstated. There are a number of UFO researchers who could have been mentioned along with their colleagues in the appendix. In the end, both space and time became constraining factors, but we wish to emphasize that this in no way lessens our appreciation of their contributions. In this spirit, the authors express their gratitude to Jan Aldrich and Richard Hall.
@@
Avi Loeb is warmly thanked for his contributions to the description of the Galileo Project and, more broadly, for his efforts to draw rigorous scientific attention to the UAP topic.
The authors thank Larry J. Hancock, Ian M. Porritt, Sean Grosvenor, Larry Cates, and Ike Okafor of SCU for providing the histograms of UAP sightings near nuclear weapon complexes (Fig. 4).
@@
Last, while it is perhaps awkward to thank one’s coauthors, KHK expresses his gratitude to them for their efforts in studying UAP (some of whom have been doing this their whole lives), for inspiring him to take on this exciting, albeit daunting, and at times, intimidating, challenge, for assisting in writing this altogether enormous and comprehensive paper, and for several years of friendship and support through what have been the most challenging and exciting years of his life. Most importantly, KHK thanks his family and especially his wife, Emily, for their support.
@@ Appendix A Prominent Past Efforts and Individuals
Four years ago, the director of Harvard College Observatory, the Administrator of NASA, the Director of the CIA and the Archbishop of Washington stood at the National Cathedral to announce that the enigma represented by UAP (UFO) reports remained an unsolved mystery. Following such statements by government and military authorities about the reality of unexplained sightings of unexplained objects in the Earth’s environment, the public and the scientific community have begun to revise their negative impressions of the subject. Naturally, they now seek reliable information on its nature. For this reason, reconstructing and presenting the history of research on the topic of UAP has become critical.
The range of phenomena known as UAP (formerly referred to as UFOs) was never accepted as a subject of interest in the Hard Sciences. Even quality studies into the material nature and behavior of reported objects have been consistently discouraged by major scientific magazines—journals such as Science, Nature or even Scientific American. Their pages continue to be denied for such studies, even when submitted by established scientists with credentials and a history of publication in other areas.
Much of the UAP literature of various quality is now found in the form of student theses, conference transcripts, and papers about witness psychology, reporting statistics, medical symptoms, and sociological interpretations of sightings, but not in formal papers about physics, engineering, or site investigation of actual cases. As for computer compilations of the many thousands of reported cases, few sources are publicly available and their reliability is debatable. Such documentation does exist, however, in the form of research studies dating back to the 1960s, but it is easy to miss because it had to be published in rare amateur magazines that have only survived in obscure Internet archives or in the confusion of online depositories. An important exception is the professionally curated and edited magazine called Flying Saucer Review, which was published in England from 1955 to 2000. The researchers who submitted their results to such an obscure publication were particularly dedicated and passionate about the subject (and no, they did not always agree).
In this section, we summarize the research efforts and works from a number of prominent PhD scientists and academics who studied UFOs relatively early on. We explicitly mentioned and described Flying Saucer Review because it is one of the only publications to which many of these individuals were able to submit their work.
@@ A.1 David Akers, Yakima Valley, Washington USA
David Akers coordinated with local residents to study anomalous light phenomena on the Yakima Indian Reservation in the vicinity of Toppenish, WA, from 1972 to the early 2000s. This activity started in 1971 and spiked in the late summer of 1972, consisting largely of nocturnal lights very near the ground or at low altitude. The initial phase of investigation lasted from August 20 to August 31, 1972, during an intense wave of sightings; many of the most interesting results were derived from this period. In this interval, nocturnal lights were reported daily in the towns of Union Gap, Harrah, Zillah, Toppenish, Wapato and Mabton. Akers used optical SLR cameras (VIS-NIR), a compass spin detector, infrared photometer, microphone and casette recorder, and energetic particle detector (Geiger-Muller counter) (see [402] Appendix B). Relying on community reports, Akers deployed his instruments where sightings were most recently reported (a ‘‘hit-and-run’’ approach) and was successful in witnessing and recording events in this way. Activity decreased dramatically after the first two days of the study.
@@
Akers acquired several photographs of what he described as ‘‘large, flickering orange objects’’ [402], although no quantitative size determination was possible. An object viewed with binoculars was described as a ‘‘sharply defined, brownish orange disc ‘‘ with a mottled surface’’. This object appeared to be flat and illuminated the ground with an orange glow. The appearance and motion of the light phenomena were highly distinctive and easily distinguished from those of conventional aircraft. On multiple occasions, attempts to acquire photographs were foiled when illumination ceased abruptly just as a camera was trained on the object. The most striking example of this occurred at a relatively close range, while an object hovered directly overhead [402].
@@
On at least one occasion, a light was recorded in a background exposure, captured when no lights were visible to the naked eye, suggesting near-infrared emission or visible band emission with a pulse duration and repetition rate below the perception threshold of human sight. Akers also noted the possibility that the objects may have been moving in a discrete or noncontinuous fashion, as suggested by the observed flickering (when observed by eye) and discontinuities in long-exposure photographs; this was not definitively distinguished from ordinary flashing and displacement while the objects are nonluminous. During the initial 11-day field campaign, none of the other instruments succeeded in recording the anomalous events. On one occasion, Akers and a colleague experienced over a minute of extremely bright and disorienting light flashes in their peripheral vision, later tentatively ascribed to strong transient magnetic field fluctuations later observed at the same location [402].
A.2 Phyllis A. Budinger, Frontier Analysis, Ltd.
Phyllis Budinger is a chemist who is one of the few scientists who has worked to study the physical traces produced during UFO encounters. She received her B.S. in chemistry in 1961 from Baldwin-Wallace University in Berea, Ohio, and her M.S. in chemistry in 1964 from Miami University in Oxford, Ohio. She began working for Sohio, which was the original Standard Oil Company that later became BP/Amoco. She accumulated 35 years of analytical chemistry experience, during which she headed a group that specialized in spectroscopy and magnetic resonance. She retired as a research scientist from BP/Amoco in 1999 [403, 404].
In 2000, Budinger founded Frontier Analysis, Ltd. as a laboratory, which has been involved in analyzing physical traces related to UFO cases, such as UFO debris [405, 406] and deposits [407, 263, 408], as in the 1971 Delphos, Kansas USA encounter [270, 403, 404], and angel hair [294], as well as chemical traces related to claims of abductions [409, 410, 411, 412], unexplained cattle mutilations [413], and crop circles.
A.3 Dr. Erol A. Faruk
@@
Dr. Erol A. Faruk was born in London, UK, in 1951. While undertaking postdoctoral organic chemical research at Nottingham University in England, he requested and received, through CUFOS, soil material that had reportedly been subjected to physical and chemical changes by an illuminated UFO observed hovering just above the ground. This event [268, 269] occurred in November 1971 on a farm in Delphos, Kansas, USA and left an eight-foot diameter ring of soil whose surface was found to glow brightly once the illuminated UFO had left the area. Subsequent field investigations showed that the ring soil had now become impervious to water down to a depth of 14 inches, such that when snow had fallen later there was no absorption by the ring, leaving only a snow-covered shape while all around was dark muddy soil.
@@
After receiving the material Faruk approached his academic supervisor (Dr. B. W. Bycroft) who, after being described the event and associated glow, became curious enough to refer him to the resident chemiluminescence expert (Dr. Frank Palmer) who became similarly intrigued by the case. Together they enabled Faruk to analyze the material while Palmer undertook recording of fluorescence spectra on samples taken as the investigation proceeded. Faruk discovered that the ring soil was impregnated by an air-sensitive organic compound that on chromatographic purification revealed it to have all the attributes necessary to generate light, a process known as oxidative chemiluminescence. He was able to use the soil data to propose a viable theory as to how the ring had actually formed, essentially confirming the witnesses’ description of the nearby presence of a hovering aerial device of unknown technology and origin [266, 267].
A.4 Dr. Pierre Guérin, Institut d’Astrophysique
@@
A.5 Dr. Richard F. Haines, NASA-Ames Research Center
@@
A.6 Paul R. Hill, NASA Langley Research Center
Paul R. Hill (1909-1990) was a distinguished American aerodynamicist who made significant contributions to mid-20th-century aerospace research. He obtained in his BS in mechanical engineering from the University of California, Berkeley in 1936, and joined the National Advisory Committee of Aeronautics (NACA), the predecessor to NASA, in 1939. He worked on various NASA aerospace projects, including aircraft design, space missions, hypersonic propulsion, and the Apollo moon missions. He was one of NASA’s chief scientists and was awarded NASA’s Exceptional Service Medal in 1969. The historian James Hansen describes Hill as one of two ‘‘key members of Langley’s early space station research.’’ [439, pp. 275–277], and as the specialist on configuration and propulsion on the HYWARDS (HYpersonic Weapon And R and D System) Program [440, p. 368].